OpenStudy (anonymous):

Last year, 20% of all claims were for weather-related damage (excluding floods), with an average weather-related damage claim being $6,500. What is the expected value (loss or gain) per weather-related claim to Acasa for every customer who chooses plan B? Plan B has a $1,200 yearly premium with a $1,000 deductible. -$1,380 -$180 $100 $1,180

OpenStudy (anonymous):

1,180

OpenStudy (anonymous):

ur wrong thankks tho

OpenStudy (anonymous):

whoops sorry

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

i need help mathmale if u could please

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

brb

OpenStudy (mathmate):

Have you worked on problems related to expected values so fr?

OpenStudy (mathmate):

Expected value is obtained by the summation: \(\Sigma~ X_i~P(X_i) \)

OpenStudy (mathmate):

Here X represents the revenue (premiums) and payout. Do not forget about the $1000 deductible, so for every claim, the payout is only $(6500-1000)=$5500

OpenStudy (mathmate):

@jcr268 I'll leave that to you to work on it. If you have questions, let me know.

OpenStudy (anonymous):

thank you . so E*X, P * (x)??

OpenStudy (mathmate):

Yes, \(E(X)=\Sigma X_i P(X_i)\) here you have two different values each for Xi and Pi. The products should be added, and it should match one of the answers.

OpenStudy (anonymous):

im confused can u put the full equation out so ik and ill solve it, just this time, so i can know what for what @Mathstudent55

OpenStudy (mathmate):

ok

OpenStudy (mathmate):

Insurance premium, X1 = 1200 (receivable from everyone) so probability P(X1)=1, and X1P(X1)=1200 Claims, X2=$6500-1000 deductible = -$5500 (pay) P(X2)=0.20 (only 20% of subscribers claim) X2*P(X2)=5500*0.2 = -1100 So \(Ex(X)=\Sigma X_i*P(X_i)=X_1P(X_1)+X_2P(X_2)=?\)